OpenStudy (anonymous):

Is this system of equations consistent, inconsistent, or dependent? 4x-3y=15 7y-35=9 1/3x A. consistent B. inconsistent C. dependent D. inconsistent and dependent E. none of these

OpenStudy (anonymous):

@DanJS can you help me? i'm leaning more towards D?

OpenStudy (danjs):

What are the definitions of each of those cases, that you know

OpenStudy (danjs):

Here is what i think: Inconsistent = no solutions to the system, no point (x,y) that satisfies both equations. They are Parallel lines. Dependent = There are infinitely many solutions to the system. This happens if both equations are the same line, they overlap at an infinite amount of points (x,y) Consistent = the lines cross at one point, and there is a unique solution (x,y) that satisfies the equations.

OpenStudy (danjs):

That sound about good?

OpenStudy (anonymous):

yes, those are basically the same definition to what i have written.

OpenStudy (danjs):

Just by those responses, without considering the equations, i think we can rule out D, i dont think it can be dependent and inconsistent at the same time can it?

OpenStudy (danjs):

And it has to be one of those 3 cases i think, i cant think of a system that is not one of the three, so that rules out E. None of the above

OpenStudy (danjs):

So there are Infinitely many solutions, No solutions, Or , 1 Unique Solution. In other words, Parallel Lines, The Same Line Overlapping, or Lines that cross Once... in a 2-D space, we aren't considering Skew lines here

OpenStudy (danjs):

So, lets take a look at the equations now

OpenStudy (danjs):

4x - 3y = 15 7y - 35 = 9 1/3x

OpenStudy (danjs):

is that 9 and 1/3 x?

OpenStudy (anonymous):

yes sorry i wasnt sure how to type that mixed fraction.

OpenStudy (danjs):

that is right, just wondering if it is a mixed number for sure

OpenStudy (anonymous):

\[9 \frac{ 1 }{ 3 }x\]

OpenStudy (danjs):

The easiest way to tell which case you have , is to try and solve the system. There are other ways that are quicker , i can tell you after trying to solve them first

OpenStudy (anonymous):

okay let me try and solve it quickly

OpenStudy (danjs):

turn that mixed number to an improper fraction first, then multiply the whole equation by the denominator to get rid of the fraction

OpenStudy (danjs):

7y - 35 = 9 1/3x 7y - (28/3)x = 35 21y - 28x = 105 ----------------------- so the two equations w/o fractions are 4x - 3y = 15 -28x + 21y = 105

OpenStudy (anonymous):

ok thats what i just got as well.

OpenStudy (danjs):

before solving, let me show you this neat trick

OpenStudy (anonymous):

alright :)

OpenStudy (danjs):

well, actually no, try to solve first

OpenStudy (anonymous):

okay one second.

OpenStudy (anonymous):

i'm solving for x and y correct?

OpenStudy (danjs):

right, you could solve one for x or y, then substitute that into the second equation

OpenStudy (danjs):

4x - 3y = 15 -28x + 21y = 105

OpenStudy (anonymous):

alright would x be 15/4 and y -5 ?

OpenStudy (danjs):

hmm, you must have gone wrong someplace... ill start going through, lets try to solve each one for y = mx + b first ok?

OpenStudy (anonymous):

okay

OpenStudy (danjs):

4x - 3y = 15 3y = 4x - 15 ------------> y = (4/3)x - 15 -28x + 21y = 105 21y = 28x + 105 ----------> y = (4/3)x + 15 ------------------------------------------- The two equations are \[y = \frac{ 4 }{ 3 }x - 15~~~and~~~y=\frac{ 4 }{ 3}x+15\]

OpenStudy (danjs):

Any part of that you don't see how to do?

OpenStudy (anonymous):

okay yeah that's what i was getting, except \[\frac{ 4 }{ 3 }x-5\]

OpenStudy (anonymous):

so that's where i got the -5. i see what you did though

OpenStudy (danjs):

ok, this is another way to find the solution,... now that you have both in the form y = mx+b set y1 = y2 and solve for x , we do that because we are looking for a point (x,y) that will work in both equations. So Y in the first equation should equal Y in the second equation. Y = Y, replace with mx+b = mx+b and solve for x

OpenStudy (danjs):

\[\frac{ 4 }{ 3}x + 15 = \frac{ 4 }{ 3 }x - 15\]

OpenStudy (danjs):

what do you get for x when you solve that?

OpenStudy (anonymous):

lets see. what if i don't get anything?

OpenStudy (danjs):

take off 4/3 x on both sides, you are left with 15 = -15 right?

OpenStudy (danjs):

That is a false statement. This means that there is NO SOLUTION (x,y) that will work in both equations.

OpenStudy (anonymous):

you're right it would bring it down to 15=-15

OpenStudy (danjs):

From the two lines in y = mx+b form... y = (4/3)x + 15 y = (4/3)x - 15 what do you notice about the slopes (m), and the Y intercepts (b)

OpenStudy (danjs):



OpenStudy (danjs):

There are no solutions to the two equations. The solutions are where the two lines intersect (x,y). Here they are parallel, They have the same slope (4/3) but different y-intercpets +15 and -15

OpenStudy (danjs):

That is why when you tried to solve it, you got a nonsense answer like -15 = 15. That means there are no solutions, and the lines are parallel.

OpenStudy (danjs):

From the initial post i did with the definitions, Which one is this then? Inconsistent = no solutions to the system, no point (x,y) that satisfies both equations. They are Parallel lines. Dependent = There are infinitely many solutions to the system. This happens if both equations are the same line, they overlap at an infinite amount of points (x,y) Consistent = the lines cross at one point, and there is a unique solution (x,y) that satisfies the equations.

OpenStudy (danjs):

I will show you a faster way to do these problems now if you would like.

OpenStudy (anonymous):

so it would solely be inconsistent then.

OpenStudy (danjs):

yep, inconsistent = no solutions -15 = 15 FALSE

OpenStudy (danjs):

Remember it like, if you get a dumb answer like 10 = 25 , obviously false, then another word for that is it is inconsistent.

OpenStudy (danjs):

Here is a fast way to tell if lines are either Parallel, or the Same Line (Inconsistent or dependent)..... 4x - 3y = 15 -28x + 21y = 105 Multiply the first equation by a constant number (-7), the system is then... -28x + 21y = -105 -28x + 21y = 105 Notice by multiplying the first equation by a constant (this doesnt change the equation at all if you do it to every term), the left hand side becomes identical for the 2

OpenStudy (danjs):

If you can multiply an equation by a constant, and the equation becomes the same left hand side, the lines are parallel.

OpenStudy (danjs):

We say , lines are parallel if one is a scalar multiple, in this case (-7) of the other line.

OpenStudy (danjs):

If you multiply through , and both sides of the equation becomes the same, then obviously you have the same line and they are dependent. For Example, Pretend the 15 in the first was a -15 4x - 3y = -15 -28x + 21y = 105 Now multiply the first equation again by (-7) and you get... -28x + 21y = 105 -28x + 21y = 105 Same exact equation, same line, dependent, infinite solutions

OpenStudy (danjs):

that is just a quicker way to tell real fast if you have parallel lines, or the same line written in 2 different forms.

OpenStudy (danjs):

Here is a little page i found to review if needed, going over each type of system... http://www.algebra.com/algebra/homework/coordinate/Types-of-systems-inconsistent-dependent-independent.lesson

OpenStudy (danjs):

if you have another one you want to run through for practice, tag me in